OpenStudy (anonymous):

If a circle has an are of 16 pi which cannot be the length of a line segment with both endpoints on the circle? 1 4 6 8 9

OpenStudy (konradzuse):

Please help the user find the answer, don't just give it out.

OpenStudy (anonymous):

Let's work backwards here. You have the area, so now you must find the radius. Let's use A=πr^2 and plug in 16π for the area: 16π=πr^2 Now you see there is pi on each side so you can divide each side by pi, giving you: 16=r^2 From here, you can determine that the radius is 4. Now the longest line that can have both endpoints in a circle is it's diameter, which is equal to 2r. This means it is 8, so every line segment 8 units or less can have both endpoints touch the edge of the circle, which means 9 is the only answer which cannot.

OpenStudy (anonymous):

Still. Not everyone understands this as much as some tutors on here. It is best to go into detail and make sure they understand why the answer is as it is.

OpenStudy (anonymous):

thank you!!!